If 8men 8hrs per day works for 8days get 45/- then how many men required if the work is 5hrs per day for10days they get 60/-?

Grade:

1 Answers

Komal
8 years ago
Let x be the number of men.
Total work done = 8*8*8 got Rs 45.
Rs 1 work done = (8*8*8)/45.......(i)
Same way, Total work1 = (5*10*x) got Rs 60
Rs 1 work done = (5*10*x)/60.......(ii)
Now from equation (i) and (ii), we get
(8*8*8)/45 = (10*5*x)/60